# Recent Advances in OOD Detection: Problems and Approaches

### v2

@@ 4.2. Test-Time Adaptive OOD Detection Approaches
Overview. Test-time adaptive methods, which employ a classifier trained on the training set, strive to utilize test data, either the complete test set or a series of unlabeled mini-batches, to enhance OOD detection performance through model adaptation. Test-time adaptive approaches are based on the theoretical insight  (Fang et al., 2022) that detecting OOD samples using only ID samples without any additional knowledge is impossible. These methods can be divided into two categories according to whether the model be modified during the testing time: Model-optimization-based and Model-optimization-free. Both of them undergo a post-training phase, during which the trained model can be adapted, regardless of whether it is updated.
@@
A line of these methods (Katz-Samuels et al., 2022; Du et al., 2024) advocate for harnessing a combination of unlabeled ID and OOD data, termed ”wild data”, which is plentiful and readily available in the real-world scenario, to OOD detection. The motivation of WOODS  (Katz-Samuels et al., 2022) is to clean the wild data to get reliable OOD candidates, and then the model regularization can be performed with the knowledge of them. Afterwards, to understand the role of wild data in OOD detection, SAL  (Du et al., 2024) explains how they help OOD detection from the lens of separability and learnability. Notably, wild data is not entirely equivalent to test OOD data if wild OOD data is not from the test datasets, inevitably leading to incorporating unintended information into the model. In addition, the extra training requirement in WOODS, thought necessary, incurs significant costs and is generally undesirable.
@@
Another line of model-optimization-based approaches draws inspiration from the semi-supervised-learning (SSL) techniques  (Zhu, 2005), aiming for a more efficient, lightweight training process during the post-training phase. Pseudo-labeling  (Arazo et al., 2020) is a simple but effective way to label the test data, which enhances learning from test unlabeled data. The method proposed by  Yang et al. (2023a), termed AUTO, employs only pseudo-OOD data to refine the model. The role of pseudo-ID data in AUTO is to mitigate catastrophic forgetting with a semantically-consistent objective, thereby maintaining the accuracy of ID classification. In contrast, ATTA  (Gao et al., 2023b) and SODA  (Geng et al., 2023) harness both pseudo-ID and pseudo-OOD data to refine the trained model. SODA employs a dual-loss approach to tackle pseudo-ID and pseudo-OOD data simultaneously, while ATTA distinguishes them with different weighting techniques.
@@
Model-optimization-free. Modifying the original trained model is infeasible in certain security-sensitive scenarios. Therefore, methods enabling test-time adaptation without requiring model updates, termed ”model-optimization-free” (MOF) techniques, are increasingly garnering interest. These approaches enhance the utilization of test data by either memorizing it or incorporating additional modules on top of the original model.
@@
Both ETLT  (Fan et al., 2022) and GOODAT  (Wang et al., 2024) retain the integrity of the original model by training an add-on module to adjust the OOD score, rather than altering the original model itself. ETLT observes a linear correlation between the feature representation and the OOD score of a test input. In other words, for a given image, the pair consisting of its feature and OOD score (feature, OOD score) exhibits a linear correlation. Furthermore, the aforementioned pairs of ID and OOD data is linearly separable. Based on the observation, ETLT proposes to learn a linear regression module trained from the (feature, OOD score) pair. The availability of complete test datasets may not always be feasible, therefore,  Fan et al. (2022) also provide an online variant to ensure safer deployment. Similarly, GOODAT develops an add-on named graph masker, designed specifically for graph data. It integrates GIB-boosted losses and employs it as the metric for OOD scoring. In contrast, AdaOOD  (Zhang et al., 2023b) avoids any additional training burden through a non-parametric k-nearest neighbours approach. The core principle behind AdaOOD is the maintenance of a memory bank, similar to the approach taken by AUTO.
@@
Online v.s. offline. Most post-hoc methods  (Liang et al., 2018b; Fan et al., 2022) have traditionally emphasized the offline scenario, where OOD detectors remain static and fixed after deployment. In contrast, the majority of test-time methods (Zhang et al., 2023b; Gao et al., 2023b) adopt the online scenario to obtain the decision boundary dynamically, minimizing the risk of incorrect OOD predictions at each time step.
@@
More challenging scenario. In the context of test-time OOD detection scenarios, some scholars have proposed more challenging configurations that demand a higher level of capability from the models. MOL (Wu et al., 2023) introduces a more realistic problem scenario, namely Continuous Adaptive Out-of-Distribution (CAOOD) detection, aimed at addressing the challenge of constantly changing ID and OOD distributions in the real world. The meta-learning approach is employed to swiftly adapt models in response to the complexities encountered in various scenarios in MOL.
Figure 9. Overview and trends in LPM-based OOD Detection Methods.
@@ 5. Problem: Large Pre-trained Model-based OOD Detection
Large pre-trained models have showcased remarkable performance in numerous downstream ID classification tasks, but their potential in OOD detection tasks remains a less-explored area. Recent research  (Vaze et al., 2022) highlights a correlation between higher ID classification accuracy and better OOD detection performance. Consequently, large pre-trained model-based OOD detection problem comes naturally. In recent years, large pre-trained models of various types, including single-modal (ViT (Dosovitskiy et al., 2021), BERT (Devlin et al., 2019), Diffusion (Ho et al., 2020)), visual language models (VLMs) (CLIP (Radford et al., 2021), multi-modal Diffusion (Saharia et al., 2022) ALIGN (Jia et al., 2021), ), and large language models (LLMs) (GPT3 (Brown et al., 2020)), have been increasingly utilized for OOD detection tasks, as shown in Fig. 9. Leveraging the powerful representational capabilities of large pre-trained models has further relaxed the constraints of OOD detection tasks, leading to a focus on more challenging and realistic scenarios, which has emerged as a new hotspot. Given the number of ID shots exposed to the large pre-trained model, large pre-trained model-based OOD detection can be classified into Zero-shot, Few-shot, and Full-shot OOD detection, as shown in Fig. 10. The performance evaluations of several relevant competitive methods are summarized in Table  2 to provide an understanding of the performance level of OOD detection in this area.
@@ 5.1. Zero-shot OOD Detection Approaches
Overview. Given the large pre-trained model and ID class names, we undertake the same task as the OOD detection, precisely detecting OOD data to abstain from prediction and accurately classifying ID data. Note that we do not need to have access to ID images and rely solely on textual category knowledge. It’s important to clarify that “ID” here refers to the ID of the specific downstream task, not the dataset ID data during pre-training.
VLMs-based. Existing research on zero-shot OOD detection using VLMs can be categorized into two main approaches based on which VLMs are utilized as a basic backbone. Diffusion-based. Traditional generative approaches, such as diffusion models  (Ho et al., 2020), capture the ID distribution effectively. Subsequently, these methods are used to identify OOD samples by assessing the likelihood that a given test input is derived from an OOD source. In a recent study in RONIN  (Nguyen et al., 2024), the diffusion model has been applied to achieve OOD object detection, which utilizes the diffusion model to generate ID inpainting, while the CLIP model is employed to calculate the similarity score. CLIP-based. Given the remarkable proficiency of CLIP in correlating texts and images, a substantial number of researchers have ventured into zero-shot OOD detection leveraging CLIP. Typically, in the zero-shot OOD detection setup, add-ons are added to the pre-trained model to better adapt it to the OOD detection task. ZOC  (Esmaeilpour et al., 2022) achieves zeto-shot OOD detection by training an image description generator on a large image captioning dataset  (Lin et al., 2014), enabling the model to generate candidate unseen labels. It’s essential to note that ZOC treats CLIP merely as a feature extractor and doesn’t impart OOD detection capabilities to CLIP itself. The subsequent work CLIPN  (Wang et al., 2023c) empowers CLIP with the ability to say “no” by a “no-prompt” encoder. Other approaches alleviate the need for additional training and focus on enhancing the performance of post-hoc zero-shot OOD detection. A simple baseline of clip-based zero-shot OOD detection is to use the normalized text-image similarity as an OOD score.  Ming et al. (2022b) further substitute the score with the maximum concept matching (MCM) score and offer a comprehensive theoretical explanation of this modification. Moreover, NegLabel  (Anonymous, 2024), LAPT (Zhang et al., 2024) and CLIPScope (Fu et al., 2024) enriches the text information through a predetermined corpus. Concretely, NegLabel selects negative words (labels) from the corpus, which are dissimilar to ID labels, to enhance OOD detection performance. Futhermore, CLIPScope uses those labels to revise the original score by applying Bayesian rules. In contrast, LAPT leverages text-to-image generation models or image retrieval models to obtain images corresponding to the given ID and Negative ID labels, followed by prompt tuning. It is noteworthy that ZOC, CLIPN, NegLabel and CLIPScope each devise unique OOD scores for integrating into their respective models. Conversely, the MCM score stands out for its generalizability across different models. The calculation of MCM is as follows:
(6) $S_{MCM}(x^{\prime};\mathcal{Y}_{in},T,\tau)=\max_{i}\frac{e^{s_{i}(x^{\prime})/\tau}}{\sum_{j=1}^{K}e^{s_{j}(x^{\prime})/\tau}},$
where $x^{\prime}$ is the test input image; $\mathcal{Y}_{in}=\{y_{1},y_{2},\ldots,y_{K}\}$ represents the label set; $y_{i}$ denotes the $i$-th label in the label set; $T(t_{i})$ corresponds to the concept vector for the text prompt $t_{i}$; $s_{i}(x^{\prime})$ is the cosine similarity score between the image feature $I(x^{\prime})$ and the concept vector $T(t_{i})$, computed as $s_{i}(x^{\prime})=\frac{I(x^{\prime})\cdot T(t_{i})}{\|I(x^{\prime})\|\|T(t_{i})\|}$; $e^{s_{i}(x^{\prime})/\tau}$ is the exponentiated similarity score normalized by the temperature parameter $\tau$; $K$ is the number of labels in the label set; $\tau$ is the temperature parameter of the softmax function; and $S_{MCM}$ is the maximum concept matching score. Some works have explored variations of zero-shot OOD detection or more challenging scenarios, such as detecting ID objects (Miyai et al., 2023c) or handling labels with noise (Ding and Pang, 2024), which will be discussed in the Section.7.
LLM-based. With the flourishing development of LLMs, new opportunities have emerged in the field of OOD detection with the application of LLMs. The strength of LLMs resides in their extensive world knowledge, enabling them to furnish comprehensive details about ID labels. Nevertheless, the issue of generating false or misleading information, known as hallucinations, poses a significant challenge when employing LLMs for OOD detection tasks. Existing methods (Huang et al., 2024; Dai et al., 2023) are based on an observation: two categories of images may have very similar visual features but different properties in the semantic space.  Dai et al. (2023) introduce a consistency-based calibration method with an object detector to mitigate hallucinations and utilize LLMs’ world knowledge for describing ID classes. However,  Huang et al. (2024) propose ODPC to generate peer-class based on ID labels and OOD samples within text and image modality. It should be noted that ODPC requires ID images for training MLP, indicating that it is not a zero-shot method. Regarding the OOD score,  Dai et al. (2023) adopt MSP, however, ODPC uses a KNN-based score. Later, a study  (Liu et al., 2024b) on LLMs was proposed to study three problems with LLMs doing OOD detection: the propensity of LLMs for near-OOD and far-OOD, the effect of different fine-tuning methods, and the choice of OOD score.  Liu et al. (2024b) find that LLMs are natural far-OOD detectors, generative fine-tuning is better than discriminative fine-tuning, and cosine distance is sufficient as OOD score because the embedding space of LLMs exhibits anisotropy.
@@
Remark. As mentioned above, “zero-shot” here refers to no exposure to ID images and only access to ID labels. However,  Fort et al. (2021) argue that in deployment scenarios, labels for OOD are either readily available. They propose a variant of the baseline approach that incorporates the names of OOD classes as candidate labels, resulting in improved performance. However, access to OOD labels is rare in real-life scenarios and seems less reasonable.
@@
Table 2. Performance evaluation of some competitive methods using the ImageNet-1K dataset as the ID dataset and iNaturalist, SUN, Places, and Textures as OOD datasets. The best result is emphasized in bold.

### v3

@@ 4.2. Test-Time Adaptive OOD Detection Approaches
Overview. Test-time adaptive methods, which employ a classifier trained on the training set, strive to utilize unlabeled test data, either the complete test set or a series of mini-batches, to enhance OOD detection performance through model adaptation. Test-time adaptive approaches are based on the theoretical insight  (Fang et al., 2022) that detecting OOD samples using only ID samples without any additional knowledge can be challenging and potentially limited. These methods can be divided into two categories according to whether the model be modified during the testing time: Model-optimization-based and Model-optimization-free. Both of them undergo a post-training phase, during which the trained model can be adapted, regardless of whether it is updated.
@@
A line of these methods (Katz-Samuels et al., 2022; Du et al., 2024a) advocate for harnessing a combination of unlabeled ID and OOD data, termed ”wild data”, which is plentiful and readily available in the real-world scenario, to OOD detection. The motivation of WOODS  (Katz-Samuels et al., 2022) is to clean the wild data to get reliable OOD candidates, and then the model regularization can be performed with the knowledge of them. Afterwards, to understand the role of wild data in OOD detection, SAL  (Du et al., 2024a) explains how they help OOD detection from the lens of separability and learnability. Notably, wild data is not entirely equivalent to test OOD data if wild OOD data is not from the test datasets, inevitably leading to incorporating unintended information into the model. In addition, the extra training requirement in WOODS, thought necessary, incurs significant costs and is generally undesirable.
@@
Another line of model-optimization-based approaches draws inspiration from the semi-supervised-learning (SSL) techniques  (Zhu, 2005), aiming for a more efficient, lightweight training process during the post-training phase. Pseudo-labeling  (Arazo et al., 2020) is a simple but effective way to label the test data, which enhances learning from test unlabeled data. The method proposed by  Yang et al. (2023a), termed AUTO, employs only pseudo-OOD data to refine the model. The role of pseudo-ID data in AUTO is to mitigate catastrophic forgetting with a semantically-consistent objective, thereby maintaining the accuracy of ID classification. In contrast, ATTA  (Gao et al., 2023a) and SODA  (Geng et al., 2023) harness both pseudo-ID and pseudo-OOD data to refine the trained model. SODA employs a dual-loss approach to tackle pseudo-ID and pseudo-OOD data simultaneously, while ATTA distinguishes them with different weighting techniques.
@@
Model-optimization-free. Modifying the original trained model is infeasible in certain security-sensitive scenarios. Therefore, methods enabling test-time adaptation without requiring model updates, termed ”model-optimization-free” techniques, are increasingly garnering interest. These approaches enhance the utilization of test data by either memorizing it or incorporating additional modules on top of the original model.
@@
Both ETLT  (Fan et al., 2022) and GOODAT  (Wang et al., 2024a) retain the integrity of the original model by training an add-on module to adjust the OOD score, rather than altering the original model itself. ETLT observes a linear correlation between the feature representation and the OOD score of a test input. In other words, for a given image, the pair consisting of its feature and OOD score (feature, OOD score) exhibits a linear correlation. Furthermore, the aforementioned pairs of ID and OOD data is linearly separable. Based on the observation, ETLT proposes to learn a linear regression module trained from the (feature, OOD score) pair. The availability of complete test datasets may not always be feasible, therefore,  Fan et al. (2022) also provide an online variant to ensure safer deployment. Similarly, GOODAT develops an add-on named graph masker, designed specifically for graph data. It integrates GIB-boosted losses and employs it as the metric for OOD scoring. In contrast, AdaOOD  (Zhang et al., 2023b) avoids any additional training burden through a non-parametric k-nearest neighbours approach. Similarly, OODD (Yang et al., 2025) constructs a dynamic OOD dictionary, managed by a priority queue, which continuously updates with feature representations of test samples based on their OOD scores. The core principle behind AdaOOD and OODD is the maintenance of a memory bank, similar to the approach taken by AUTO.
@@
Online v.s. Offline. Most post-hoc methods  (Liang et al., 2018b; Fan et al., 2022) have traditionally emphasized the offline scenario, where OOD detectors remain static and fixed after deployment. In contrast, the majority of test-time methods (Zhang et al., 2023b; Gao et al., 2023a) adopt the online scenario to obtain the decision boundary dynamically, minimizing the risk of incorrect OOD predictions at each time step.
@@
More challenging scenario. In the context of test-time OOD detection scenarios, some researchers have proposed more challenging configurations that demand a higher level of capability from the models. MOL (Wu et al., 2023a) introduces a more realistic problem scenario, namely Continuous Adaptive Out-of-Distribution (CAOOD) detection, aimed at addressing the challenge of constantly changing ID and OOD distributions in the real world. The meta-learning approach is employed to swiftly adapt models in response to the complexities encountered in various scenarios in MOL.
Figure 5. Overview and trends in LPM-based OOD Detection Methods. (a) presents statistical data on influential research works in recent years, which involves papers indexed in prominent journals or conferences in the field. In the bar chart, the lower and upper parts represent the total number of papers in the first and second halves of each year, respectively. (b) shows a trend pie chart illustrating the proportion of keyword categories related to the research works in that year.
@@ 5. Problem: Large Pre-trained Model-based OOD Detection
Large pre-trained models have showcased remarkable performance in numerous downstream ID classification tasks, but their potential in OOD detection tasks remains a less-explored area. Recent research  (Vaze et al., 2022) highlights a correlation between higher ID classification accuracy and better OOD detection performance. Consequently, large pre-trained model-based (LPM-based) OOD detection problem comes naturally. In recent years, large pre-trained models of various types, including single-modal models (ViT (Dosovitskiy et al., 2021), BERT (Devlin et al., 2019), Diffusion (Ho et al., 2020)), visual language models (VLMs) (CLIP (Radford et al., 2021), multi-modal Diffusion (Saharia et al., 2022), ALIGN (Jia et al., 2021), ), and large language models (LLMs) (GPT3 (Brown et al., 2020)), have been increasingly utilized for OOD detection tasks, as shown in Fig. 5. Based on the trend chart, it can be anticipated that LPM-based OOD Detection will continue to be a key area of focus in the future. Leveraging the powerful representational capabilities of large pre-trained models has further relaxed the constraints of OOD detection tasks, leading to a focus on more challenging and realistic scenarios, which has emerged as a new hotspot. Given the number of ID shots exposed to the large pre-trained model, OOD detection in this section can be classified into Zero-shot, Few-shot, and Full-shot OOD detection, as shown in Fig. 6. The performance evaluations of several relevant competitive methods are summarized in Table  3 to provide an understanding of the performance level of OOD detection in this area.
@@ 5.1. Zero-shot OOD Detection Approaches
Overview. Given the large pre-trained model and ID class names, we undertake the same task as the OOD detection, precisely detecting OOD data to abstain from prediction and accurately classifying ID data. Note that we rely solely on textual category knowledge, without the need for access to ID images. It’s important to clarify that “ID” here refers to the ID of the specific downstream task, not the dataset ID data during pre-training. Existing research on zero-shot OOD detection using vision-language models (VLMs-based) can be broadly categorized into two main approaches, depending on the underlying models employed: diffusion-based and CLIP-based methods.
Diffusion-based. Recent generative approaches, such as diffusion models (Ho et al., 2020), effectively capture the ID data by modeling its underlying distribution. These models typically detect OOD samples by evaluating the likelihood of a test input belonging to the modeled ID distribution. For instance, RONIN (Nguyen et al., 2024) leverages a diffusion model for ID inpainting, combined with the CLIP model to compute similarity scores for OOD detection. However, most existing works focus more extensively on CLIP-based methods due to their superior zero-shot capabilities and flexibility.
CLIP-based. In contrast to diffusion-based approaches, CLIP-based methods have received significant attention for zero-shot OOD detection. These methods exploit the powerful vision-language alignment capability of CLIP to perform OOD detection without requiring additional training on OOD data. In the following, we focus on the development and evolution of CLIP-based zero-shot OOD detection methods, discussing both approaches that require extra training and post-hoc methods, as well as recent advances leveraging textual information.
A representative approach, ZOC (Esmaeilpour et al., 2022), introduces an image description generator trained on large-scale image captioning datasets (Lin et al., 2014), enabling the generation of candidate unseen labels for OOD detection. Notably, ZOC treats CLIP primarily as a feature extractor and does not inherently endow it with OOD discrimination capabilities. Building upon this, CLIPN (Wang et al., 2023c) enhances CLIP’s expressiveness by incorporating a “no-prompt” encoder, thereby empowering CLIP to explicitly reject unfamiliar inputs—a critical step toward robust OOD detection in open-world scenarios. Beyond methods that require further training, a parallel line of research investigates post-hoc zero-shot OOD detection, aiming to maximize CLIP’s potential without additional fine-tuning. The most straightforward baseline in this category computes the normalized text-image similarity as the OOD score. To improve upon this, MCM (Ming et al., 2022b) replaces the similarity score with a Maximum Concept Matching (MCM) score, providing a theoretically grounded and empirically robust alternative. The MCM score is defined as:
(6) $S_{MCM}(x^{\prime};Y_{in},T,\tau)=\max_{i}\frac{e^{s_{i}(x^{\prime})/\tau}}{\sum_{j=1}^{K}e^{s_{j}(x^{\prime})/\tau}}.$
where $x’$ ’ denotes the test image, $Y_{in}$ is the set of in-distribution labels, $T(t_{i})$ is the text embedding for prompt $t_{i}$ , $s_{i}(x’)$ ’ ) is the cosine similarity between the image and text features, and $\tau$ is the softmax temperature.
Recent works have further enriched the textual context used by CLIP to improve OOD detection. For example, NegLabel (Anonymous, 2024), LAPT (Zhang et al., 2024a), CSP (Chen et al., 2024b), and CLIPScope (Fu et al., 2024) all leverage curated corpora to augment the set of candidate labels. Specifically, NegLabel selects negative labels from a corpus that are semantically dissimilar to ID classes, thereby sharpening the model’s ability to distinguish OOD samples. CLIPScope (Fu et al., 2024) goes a step further by integrating these negative labels into a Bayesian scoring framework to revise the original OOD score. LAPT, in contrast, utilizes text-to-image generation or image retrieval models to obtain visual exemplars for both ID and negative labels, followed by prompt tuning to optimize OOD discrimination. It is worth emphasizing that methods such as ZOC (Esmaeilpour et al., 2022), CLIPN (Wang et al., 2023c), NegLabel (Anonymous, 2024), and CLIPScope (Fu et al., 2024) each propose distinct OOD scoring mechanisms tailored to their respective architectures. In contrast, the MCM score stands out for its generalizability and can be seamlessly integrated into diverse CLIP-based frameworks.
Recent advances in zero-shot OOD detection have moved beyond static negative sampling by leveraging more adaptive strategies within the joint vision-language space. For example, AdaNeg (Zhang and Zhang, 2024) dynamically generates negative proxies at test time to better match the semantics of OOD samples, while AdaND (Cao et al., 2025) focuses on mitigating the impact of noisy samples by decoupling noise detection from classification using CLIP-based pseudo-labels. In contrast, SimLabel (Zou et al., 2025) distinguishes OOD inputs by analyzing the consistency of image-label similarities among semantically related classes, automatically constructing similar class groups via hierarchical structures and large language models. These approaches thus differ in their strategies for negative sampling, noise handling, and semantic grouping, reflecting the diverse directions in recent OOD detection research.
Complementing these proxy- and consistency-based advances, several works have proposed to enhance OOD detection by expanding the diversity and expressiveness of negative or outlier labels. OLE (Ding and Pang, 2024) exposes CLIP to a large set of outlier labels, selected from external corpora and clustered into prototypes, to better model the outlier space. Furthermore, OLE augments these prototypes through interpolation with ID representations, effectively filling the semantic gap between ID and OOD regions in the embedding space. Similarly, recent work on prompt learning for near-OOD detection (Jung et al., 2024) introduces novel scoring functions, such as the Empty-Class Score, which measures the alignment between an image and the context vector excluding any class name, and relative scores that subtract this alignment from standard logit-based OOD scores. This simple post-hoc adjustment provides substantial gains, especially for challenging near-OOD cases, without requiring retraining or architectural changes.
Beyond label- and score-centric approaches, matrix approximation techniques have also been explored. SeTAR (Li et al., 2024b) applies selective low-rank approximation to CLIP’s weight matrices, aiming to retain the most informative components for OOD discrimination while filtering out spurious correlations. In a different vein, OT-DETECTOR (Liu et al., 2025a) formulates OOD detection as an optimal transport problem, quantifying the semantic and distributional discrepancy between test samples and ID labels, and further refining detection by semantic-aware content refinement.
Collectively, these developments illustrate a clear methodological evolution: from static post-hoc scoring and negative label augmentation, to adaptive proxy generation, label consistency modeling, prototype-based outlier exposure, and low-rank or optimal transport-based representations. This progression not only improves OOD detection accuracy but also enhances robustness in multi-label, semantically rich, and highly dynamic deployment scenarios. Furthermore, the interplay between proxy-based, consistency-based, and low-rank approximation methods suggests promising directions for unified frameworks that leverage the strengths of each paradigm.
LLM-based. With the rapid advancement of LLMs, new opportunities have emerged in the field of OOD detection by leveraging the extensive world knowledge and powerful semantic reasoning capabilities of LLMs. The primary strength of LLMs lies in their ability to provide comprehensive and context-rich descriptions of ID labels, which can be crucial for distinguishing between ID and OOD samples. However, a significant challenge in employing LLMs for OOD detection is their tendency to generate hallucinations—false or misleading information—which can undermine the reliability of OOD identification.
Recent works have sought to address these challenges by exploring the nuanced relationship between visual and semantic spaces. For instance, existing methods (Huang et al., 2024; Dai et al., 2023) observe that visually similar image categories may possess distinct properties in the semantic space, which can be exploited for more effective OOD detection. Dai et al. (2023) introduce a consistency-based calibration method that incorporates an object detector to mitigate hallucinations, thereby leveraging the world knowledge of LLMs for more accurate ID class descriptions. In contrast, Huang et al. (2024) propose ODPC, which generates peer classes based on both ID labels and OOD samples across textual and visual modalities. Notably, ODPC requires ID images for training an MLP, indicating that it is not a zero-shot approach. In terms of OOD scoring, Dai et al. (2023) adopt the MSP, while ODPC utilizes a KNN-based scoring strategy.
Building upon these foundations, subsequent studies have systematically investigated the behavior of LLMs in OOD detection. For example, Liu et al. (2024a) analyze the propensity of LLMs to detect near-OOD versus far-OOD samples, the impact of different fine-tuning strategies, and the suitability of various OOD scoring functions. Their findings reveal that LLMs are naturally adept at detecting far-OOD samples, that generative fine-tuning outperforms discriminative fine-tuning, and that cosine distance is an effective OOD score due to the anisotropic nature of LLM embedding spaces.
More recently, LLM-based approaches such as CMA (Lee et al., 2025), COOD (Liu et al., 2024c), and EOE (Cao et al., 2024) have extended OOD detection into multi-label and concept-rich regimes. These methods utilize LLMs not only as knowledge generators, but also as proxy creators and synthetic data generators (Abbas et al., 2025), thereby enriching the semantic context available for OOD detection. The synergy between VLMs for visual grounding and LLMs for semantic reasoning has been further highlighted by works such as ReGuide (Kim et al., 2025), which propose self-guided, image-adaptive concept generation to enhance robustness. Collectively, these advancements demonstrate the growing potential of LLM-based techniques to address the complex challenges of OOD detection by effectively integrating visual and semantic information.
@@
Remark. As mentioned above, “zero-shot” here refers to no exposure to ID images and only access to ID labels. However,  Fort et al. (2021); Ding and Pang (2024) argue that in deployment, labels for OOD are either readily available. They propose a variant of the baseline approach that incorporates the names of OOD classes as candidate labels, resulting in improved performance. However, access to OOD labels is rare in real-life scenarios and seems less reasonable.
@@
Table 3. Performance evaluation of some competitive methods based on CLIP-ViT-B/16 using the ImageNet-1K dataset as the ID dataset and iNaturalist, SUN, Places, and Textures as OOD datasets. The results are cited from (Ming et al., 2022b; Chen et al., 2024b; Zhang and Zhang, 2024; Li et al., 2024c) and the best result is emphasized in bold.
@@
LAPT  (Zhang et al., 2024a) 99.63 1.16 96.01 19.12 92.01 33.01 91.06 40.32 94.68 23.40
CSP  (Chen et al., 2024b) 99.60 1.54 96.66 13.66 92.90 29.32 93.86 25.52 95.76 17.51
AdaNeg  (Zhang and Zhang, 2024) 99.71 0.59 97.44 9.50 94.55 34.34 94.93 31.27 96.66 18.92